# Solving the Inverse Problem of Electrocardiography for Cardiac Digital Twins: A Survey

### v2

@@ 6 Discussion and Future Perspectives
6.1 Surrogate Model for Efficient ECG Simulation
@@
Traditional methods for simulating ECG signals involve complex biophysical models that account for intricate cardiac electrical and anatomical properties. While these models offer high fidelity, they often require substantial computational resources and time [5]. Fast implementations of these models exist. For example, Sachetto et al. [212] presents a GPU-based implementation of the monodomain model, which is capable of speeding up simulations by a factor up to 498 when compared to a serial implementation of the model. They reported simulation times of 32 minutes per heartbeat. Other studies have proposed alternative models that can yield similar simulation outputs at a fraction of the state-of-the-art monodomain models. For example, Wallman et al. [121] proposed a graph-based Eikonal model that can simulate the activation sequence in the heart in less than 1.6 seconds [103]. Similarly, Neic et al. [213] presented a pseudo-monodomain model that can produce nearly identical simulations to a monodomain model (under sinus rhythm) but with a computational cost of 213 seconds per heartbeat [214]. Nevertheless, in the context of generating digital twins, the number of simulations required per patient can easily be in the range of hundreds of thousands [103, 138]. Thus, even the fastest mechanistic models available present a significant computational burden when designing studies involving the generation of digital twins at scale [214, 138].
@@
The emergence of neural surrogate models addresses these challenges by offering computationally efficient alternatives that balance accuracy with practicality [215, 216, 217]. These surrogate models leverage machine learning techniques to learn the intricate mapping between the physiological parameters (e.g., earliest activation sites) and expected behavior (e.g., cardiac activation sequences) [217]. For example, Bertrand et al. [217] was able to produce simulations of the activation sequence in 0.04 seconds, which was a 500-factor speed-up compared to a serial implementation of the graph-based Eikonal [121] (20 seconds per simulation). By circumventing the need for exhaustive simulations, surrogate models provide an avenue for rapid and on-demand generation of cardiac functional data (e.g., ECG signals). This is particularly valuable in scenarios where real-time simulations or large-scale analyses are necessary. However, challenges such as model interpretability, generalization across diverse populations, and the need for accurate training data warrant careful consideration. Furthermore, a thorough understanding of the trade-off between computational efficiency and simulation accuracy is vital when integrating surrogate models into clinical applications. As surrogate models continue to evolve, they hold the potential to revolutionize ECG inverse inference by enabling researchers and practitioners to efficiently explore the parameter space and improve our understanding of complex cardiac phenomena.
@@ 6.2 Multimodal Representation Learning Towards Solving the ECG Inverse Problem
The ECG inverse inference involves the fusion of information from multiple modalities, such as ECG signals, anatomical images (e.g., CT/ MRI scans), and clinical metadata. Multimodel representation learning offers a promising solution by allowing the integration of diverse data sources, leading to improved insights into the complex dynamics of cardiac electrical behavior. This can help address the issue of limited electrode coverage by inferring missing data points through learned relationships from other modalities. Furthermore, the inclusion of anatomical images can enhance the anatomical accuracy of the models, leading to improved source localization and more accurate inverse solutions. Indeed, deep neural networks have proven to be a highly effective strategy for ECG interpretation and analysis via spatiotemporal representation learning [47]. However, most works only considered ECG data for analysis and only a few works employed joint analysis with other data sources [218, 189, 219, 116]. Among these works, VAE has been widely employed to learn a latent representation of ECG and geometry [220], along with incorporating uncertainty modelling for inverse inference [116]. However, these studies merely concatenated the acquired features or utilized a shared representation, lacking a thorough explanation of how these features align and contribute to holistic data comprehension. In contrast, physics-informed models can integrate domain-specific knowledge and principles derived from the underlying physics of the cardiac system [179]. By explicitly encoding the physical interactions and dependencies between ECG and geometry, these models could offer a more transparent and interpretable way to capture the intricate interplay between these data modalities. Moreover, they could potentially enhance the generalization capabilities of the learned representations, making them more robust in scenarios with limited or noisy data. Nevertheless, the integration of physics-based constraints into machine learning models introduces additional challenges, such as learning geometrical invariance, as well as processing model complexity and the need for accurate domain knowledge. Balancing the complexity of the physical model with the capacity of the machine learning algorithm while also ensuring the interpretability of the resulting features will require careful consideration and innovative methodologies.
6.3 The Role of Cardiac Digital Twins in Advancing in Silico Clinical Trials
Cardiac digital twins are revolutionizing the landscape of in silico clinical trials, offering unprecedented opportunities to accelerate drug development [138]. By creating virtual replicas of the individual patient’s heart, cardiac digital twins enable highly personalized simulations that capture the intricate dynamics of cardiac function and response to interventions [221]. Traditional clinical trials often face limitations in recruiting participants from specific demographic or clinical groups, leading to challenges in generalizing findings to broader populations [222]. With cardiac digital twins, researchers can create virtual patient cohorts that span a wide range of demographic, clinical, and pathological characteristics. This diversity allows for a more comprehensive exploration of drug effects and treatment outcomes across different patient subgroups, enhancing the reliability and generalizability of trial results. At the same time, in silico trials significantly reduce the time and cost associated with recruiting participants, conducting experiments, and analyzing results. This accelerated pace facilitates the rapid iteration and refinement of drug candidates, ultimately expediting the journey from discovery to market availability. Moreover, in silico trials enable researchers to explore a broader range of hypotheses and scenarios than would be feasible in traditional clinical settings. This flexibility allows for a more comprehensive exploration of potential drug effects, interactions, and mechanisms of action, leading to deeper insights and more informed decision-making. In the future, in silico clinical trials hold immense potential for revolutionizing drug development and healthcare delivery. The integration of patient-specific data and virtual patient models will enable personalized drug development approaches tailored to individual characteristics and needs. To fully realize the potential of in silico clinical trials, challenges such as model validity, regulatory compliance, and data access barriers must be addressed.
@@ 7 Conclusion
Cardiac digital twins are pivotal in advancing precision medicine and improving patient outcomes in cardiovascular care by simulating and predicting personalized responses to interventions. However, current models encounter a computational bottleneck in calibrating to an individual’s physiology through ECG inverse inference, particularly in 3D heart models. This review comprehensively outlines recent progress in ECG inverse inference, encompassing algorithms, public datasets, evaluation metrics, clinical applications, and future perspectives. While machine learning, especially deep learning, holds promise in tackling this complex and ill-posed problem, significant challenges persist, alongside promising opportunities for further advancement and innovation. In the future, as these techniques gain increasing popularity and generate massive cohorts of digital twins, they are poised to revolutionize cardiovascular care through in-silico trials and personalized interventions. With the proliferation of big data, envision a future where everyone possesses a digital twin, significantly reducing the reliance on animal models and enabling real-time monitoring in hospitals.

### v4

@@ 5.1 Quantifying and Mitigating Uncertainties in ECG Inverse Inference
5.2 Surrogate Model for Efficient ECG Simulation in Cardiac Digital Twin Development
@@
The functional twinning of cardiac EP often requires extensive simulations to achieve personalized parameter optimization. Traditional methods for simulating ECG signals involve complex biophysical models that account for intricate cardiac electrical and anatomical properties. While these models offer high fidelity, they often require substantial computational resources and time [251]. Fast implementations of these models exist. For example, Sachetto et al. [252] presents a GPU-based implementation of the monodomain model, which is capable of speeding up simulations by a factor up to 498 when compared to a serial implementation of the model. They reported simulation times of 32 min per heartbeat. Other studies have proposed alternative models that can yield similar simulation outputs at a fraction of the state-of-the-art monodomain models. For example, Wallman et al. [253] proposed a graph-based Eikonal model that can simulate the activation sequence in the heart in less than 1.6 s [1]. Similarly, Neic et al. [254] introduced a reaction-Eikonal model, later implemented with a full pseudo-bidomain approach by Gillette et al. [2], which reported a computational time of 213 s per heartbeat. A simplified version using a lead field approach, without repolarization, can potentially be 10 times faster. These studies demonstrate a range of efficient ECG simulation tools, though reported times can vary significantly depending on model complexity (e.g., isotropic vs. anisotropic models), mesh resolution, and the hardware used in each approach [255, 2]. Nevertheless, in the context of generating digital twins, the number of simulations required per patient can easily be in the range of hundreds of thousands [1, 194]. Thus, even the fastest EP models available present a significant computational burden when applied at scale for digital twin generation [2, 194].
@@
The emergence of neural surrogate models addresses these challenges by offering computationally efficient alternatives that balance accuracy with practicality [119, 256, 257]. These surrogate models leverage machine learning techniques to learn the intricate mapping between the physiological parameters (e.g., EASs) and expected behavior (e.g., cardiac activation sequences) [257]. For example, Bertrand et al. [257] was able to produce simulations of the activation sequence based on U-Net in 0.04 s, which was a 500-factor speed-up compared to a serial implementation of the graph-based Eikonal. Their surrogate EP model was generic and simplified (CV was fixed) and required an additional 5h 32 min of training time, but once trained, it can be deployed with high efficiency. Instead, Jiang et al. [256] developed a personalized neural surrogate model using meta-learning, reducing simulation time to 0.24 s and personalization time to 0.032 s, with an overall training time of about 8.5 min, significantly faster than the conventional 5 min AP simulation requiring 100 simulation calls. By circumventing the need for exhaustive simulations, surrogate models provide an avenue for rapid and on-demand generation of cardiac functional data (e.g., ECG signals). This is particularly valuable in scenarios where real-time simulations or large-scale analyses are necessary. However, challenges such as model interpretability, generalization across diverse populations, and the need for accurate training data warrant careful consideration. Furthermore, a thorough understanding of the trade-off between computational efficiency and simulation accuracy is vital when integrating surrogate models into clinical applications. As surrogate models continue to evolve, they hold the potential to revolutionize CDTs by enabling researchers to efficiently explore the parameter space and improve our understanding of complex cardiac phenomena.
@@ 5.3 Multimodal Representation Learning Towards Solving the ECG Inverse Problem
The ECG inverse inference involves the fusion of information from multiple modalities, such as ECG signals, anatomical images (e.g., CT/ MRI scans), and EP parameters. Multimodal representation learning, a technique widely used in the computer vision field, holds significant potential for ECG inverse inference by enabling the seamless integration of diverse data sources [258]. Deep neural networks have proven to be a highly effective strategy for ECG interpretation and analysis via spatiotemporal representation learning [259]. However, most works only considered ECG data for analysis and only a few works employed joint analysis with other data sources [102, 147, 260, 188]. Among these multimodal works, VAE has been widely employed to learn a latent representation of ECG and geometry [261, 188], along with incorporating uncertainty modeling for inverse inference [102]. However, these studies merely concatenated the acquired features or utilized a shared representation, lacking a thorough explanation of how these features align and contribute to holistic data comprehension. In contrast, physics-informed models can integrate domain-specific knowledge and principles derived from the underlying physics of the cardiac system [121]. By explicitly encoding the physical interactions and dependencies between ECG and geometry, these models could offer a more transparent and interpretable way to capture the intricate interplay between these data modalities. Moreover, they could potentially enhance the generalization capabilities of the learned representations, making them more robust in scenarios with limited or noisy data. Nevertheless, the integration of physics-based constraints into machine learning models introduces additional challenges, such as learning geometrical invariance, as well as processing model complexity and the need for accurate domain knowledge. Balancing the complexity of the physical model with the capacity of the machine learning algorithm while also ensuring the interpretability of the resulting features will require careful consideration and innovative methodologies.
@@ 6 Conclusion
CDTs are pivotal in advancing precision medicine and improving patient outcomes in cardiovascular care by simulating and predicting personalized responses to interventions. However, current models encounter a computational bottleneck in accurately calibrating CDTs to individual physiology through ECG inverse inference, particularly in estimating EP parameters within 3D heart models. This review comprehensively outlines recent progress in ECG inverse inference, encompassing algorithms, public datasets, evaluation, clinical applications, and future perspectives. While machine learning, especially deep learning, holds promise in tackling this complex and ill-posed problem, significant challenges persist, alongside promising opportunities for further advancement and innovation. In the future, as these techniques gain increasing popularity and generate massive cohorts of digital twins, they are poised to revolutionize cardiovascular care through in-silico trials and personalized interventions. With the growing availability of big data, we can envision a future where everyone possesses a digital twin, significantly reducing the reliance on animal models and enabling real-time monitoring and precision treatment in clinical settings.